Let A be a 3 x 3 matrix such that A^2 = 2A. Prove that det(A) must be either 0 or 8.
1
Expert's answer
2020-02-17T12:18:07-0500

As per the given question,

"A^2=2A"

So, "A^2-2A=0"

"A(A-2I)=0"

So, A=0 or A=2I

As we know, "A^2=2A" is possible only for the diagonal matrix.

Here the given matrix is order of 3x3

So, det(A)=0 or det(A)=8